# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ I.1 Exponential Inner Product
$s=O\Big{(}\max\Big{\{}\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)/R)},R^{2}\Big{\}}\Big{)},$ (9)
@@
such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition I.1) of $A\in\mathbb{R}^{n\times n}$. Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
Here we consider the vector version of Lemma I.2.
Definition I.3.
@@
Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in AS [6].
Remark I.4.
@@
Lemma I.5 (Polynomial approximation).
For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in AS [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Then, for any $x,y\in[0,R]^{d}$, we have
@@
Let $n=1$. The proof follows from directly applying Lemma I.2. ∎
Using the results from AS [6] above, we can extend our results to Softmax activation.
Lemma I.6 (Weighted Softmax approximation, formal version of Lemma C.7).
Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5. Then we can approximate exponential inner product using polynomial kernel function:
$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s})$
@@
From Lemma I.5, we can use polynomial kernel to approximate the Softmax function:
$w^{\top}\exp(Xy/d)=$ $~{}\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)+O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma I.5.
@@
I.2 Algorithm Modifications
Based on Lemma I.6, we can now extend our DP algorithms to handle Softmax activation. First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time. Then, for the second term in Lemma I.6, i.e. $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm. For the first term, $-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.
Due to the decomposability of $\ell_{p}^{p}$ norm, i.e.
$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p},$
we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma E.2). We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 5 and 6) to $\ell_{2}^{2}$ distance with minor modifications.
Theorem I.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
With $\alpha$ scaled down by a factor of $2$ and all Query instead multiplied by $R^{2}/(1+\alpha/2)^{2j}$ in Lines 8 and 13 of Algorithm 6, i.e., from
  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$
to
  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R^{2}}{(1+\alpha/2)^{2j}}$.
The data structure DPTreeDistance (Algorithm 5,6) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6)
    It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}|\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}+O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$,
@@ I.2 Algorithm Modifications
The proof is similar to that of Theorem F.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$. ∎
Now we can give our result that can answer Softmax query.
Theorem I.8 (Softmax query, formal version of Theorem 4.2).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition I.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$. (Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
The DP proof is the same as the proof of Lemma G.1.
We then show the time complexity. From Lemma I.6, we know that constructing $P(\cdot)$ requires $O(r)$ time. In the first for loop of Init, the dominating time consumption is $O(nr)$. The second for loop also has a time complexity of $O(nr)$. Therefore, the total time complexity for Init is $O(nr)$. In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time. Within the for loop, it requires $O(\alpha^{-1}r\log^{2}n)$. Thus, the total time complexity for DistanceQuery is $O(\alpha^{-1}r\log^{2}n)$.
@@
The proof of the error follows from the triangle inequality by combining the errors in Lemma I.6 and Theorem I.7. ∎
I.3 Adaptive Softmax
In this section, we show how to make Algorithm 3 robust to adaptive query. We follow the same idea from Section H. We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section H. Therefore, we need to recalculate Lipschitz constant first.
Lemma I.9 (Lipschitz of weighted Softmax).
@@ I.3 Adaptive Softmax
where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact B.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$. ∎
@@
Lemma I.10 (Adaptive Softmax, formal version of Lemma C.8).
@@
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
@@
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
@@
$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
@@
The proof follows from the same idea as the proof of Lemma H.8, except that we use Theorem I.8 and the Lipschitz in Lemma I.9. ∎
Theorem I.11 (Adaptive query Softmax data structure, formal version of Theorem 4.4).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition I.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnr)$ spaces to solve weighted Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. (Algorithm 1) It takes $O(lnd)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 1) It takes $O(\alpha^{-1}ld\log^{2}n)$ time to output a number $z$ such that
@@
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma I.10. We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$. It can be shown that:
@@
I.4 Proof of Main Result

### v3

@@ G.1 Exponential Inner Product
$s=O\Big(\max\Big\{\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)/R^{2})},R^{2}\Big\}\Big),$ (9)
@@
such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition G.1) of $A\in\mathbb{R}^{n\times n}$. Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
Here we consider the vector version of Lemma G.2.
Definition G.3.
@@
Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in [4].
Remark G.4.
@@
Lemma G.5 (Polynomial approximation).
For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in [4] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Then, for any $x,y\in[0,R]^{d}$, we have
@@
Let $n=1$. The proof follows from directly applying Lemma G.2. ∎
Using the results from [4] above, we can extend our results to Softmax activation.
Lemma G.6 (Weighted Softmax approximation ).
Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5. Then we can approximate exponential inner product using polynomial kernel function:
$|-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})-w^{\top}\exp(Xy/d)|$
$=$ $~O(|w^{\top}\exp(Xy/d)\cdot\epsilon_{s}|)$
@@
From Lemma G.5, we can use a polynomial kernel to approximate the Softmax function:
$|\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)-w^{\top}\exp(Xy/d)|=$ $O(|w^{\top}\exp(Xy/d)\cdot\epsilon_{s}|).$
The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma G.5.
@@
G.2 Algorithm Modifications
Based on Lemma G.6, we can now extend our DP algorithms to handle Softmax activation. First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time. For the first term, $\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it. For the second term in Lemma G.6, i.e., $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|_{2}^{2}+\|P(y)\|_{2}^{2})$, it can be expressed as $\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-0|^{2}$ and $\frac{1}{2}\sum_{i\in[n]}w_{i}(\sum_{j\in[r]}P(y)_{j}^{2})$. The former can be computed using query 0, while the latter can be solved using the precomputed value $\sum_{i\in[n]}w_{i}$, which can be obtained from the data ${\bf 1}_{n}$ and query 0.
Thus, we only need to consider the case $p=2$ in weighted $\ell_{p}^{p}$ distance algorithms.
Now we can give our result that can answer Softmax query.
Theorem G.7 (Softmax query, formal version of Theorem 5.2).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
[…]
    • $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@ G.2 Algorithm Modifications
Let $P_{wx}:=\sum_{i\in[n]}w_{i}\|P(x_{i})\|^{2}_{2}$ and $s_{w}:=\sum_{i\in[n]}w_{i}$. Observe that $P_{wx}=\sum_{i\in[n]}w_{i}\|P(x_{i})-0\|^{2}_{2}$, meaning we can calculating $P_{wx}$ using query 0. Similarly, $s_{w}=\sum_{i\in[n]}w_{i}\|{\bf 1}_{n}-0\|^{2}_{2}$, meaning we can calculating $s_{w}$ using data ${\bf 1}_{n}$ and query 0. Thus, we compute $P_{wx},s_{w}$ in Line 16 and 19 in Algorithm 3 in this way.
From the privacy proof of Lemma E.1 and the way we choose privacy parameters, similarly we get the output process of calculating $P_{wx}$ and Value is $(\epsilon/3,\delta/3+\delta^{\prime}/2)$-DP. Also, the output process of calculating $s_{w}$ is $(\epsilon/3,\delta/3)$-DP. Then, by Fact A.8, overall process is $(\epsilon,\delta+\delta^{\prime})$-DP in Line 28 of Algorithm 3.
We then show the time complexity. From Lemma G.6, we know that constructing $P(\cdot)$ requires $O(r)$ time. In the first for loop of Init, the dominating time consumption is $O(nr)$. The second for loop also has a time complexity of $O(nr)$. Therefore, the total time complexity for Init is $O(nr)$. In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time. Within the for loop, it requires $O(r\log n)$. Thus, the total time complexity for DistanceQuery is $O(r\log n)$.
@@
The proof of the error follows from the triangle inequality by combining the errors in Lemma G.6 and Theorem E.3. Here, we omit the constant factors of 2 and 3 used for the privacy guarantee in Algorithm 3, incorporating it into the big-$O$ notation for the error analysis. To be more specific, in Line 28 of Algorithm 3, we have 3 terms to bound the error, namely $P_{wx},s_{w}\|P(y)\|_{2}^{2}$ and Value. From Lemma G.6, the first source of error comes from the approximation error introduced by polynomial kernel method, i.e.,
$~|w^{\top}\exp(Xy/d)-\frac{1}{2}(\underbrace{\sum_{i\in[n]}w_{i}\|P(x_{i})\|^{2}_{2}}_{P_{wx}}+\underbrace{\sum_{i\in[n]}w_{i}}_{s_{w}}\|P(y)\|^{2}_{2}-\underbrace{\sum_{i\in[n]}w_{i}\|P(x_{i})-P(y)\|_{2}^{2}}_{\mathrm{Value}})|$
$=$ $~O(|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|).$
Then, the second source of error comes from the DP noises in Theorem E.3, where we use Algorithm 4 to compute the three terms.
The two terms $P_{wx}$ and Value have additive error $O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$ (Theorem E.3) due to to the way we choose the DP parameters, the appl==ication of advanced composition (Theorem A.10), and the transformation of the value range from $[0,R]$ to $0,\Gamma_{R,s}]$ by the polynomial kernel. See more details in the proof of Lemma E.2.
As for the term $s_{w}\|P(y)\|_{2}^{2}$, the addtive error of $s_{w}$ is $O(\epsilon^{-1}R_{w}\log^{3/2}n)$. But since $\|P(y)\|_{2}^{2}\leq r\Gamma_{R,s}^{2}$, we have the addtive error is $O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\log^{3/2}n)$ which is smaller than other two terms. We ignore the constant 3 introduced by summing three terms by triangle inequality of absolute function, i.e., $|-t_{1}+t_{2}+t_{3}|\leq|t_{1}|+|t_{2}|+|t_{3}|$.
Finally, summing the two sources of error by triangle inequality, we finish the proof. ∎
G.3 Adaptive Softmax
In this section, we show how to make Algorithm 3 robust to adaptive query. We follow the same idea from Section F. We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$ distance in Section F. Therefore, we need to recalculate Lipschitz constant first.
Lemma G.8 (Lipschitz of weighted Softmax).
@@ G.3 Adaptive Softmax
where the first step follows from definition of Z(y),Z(y~)Z(y),Z($\widetilde${y}), the second step follows from triangular inequality, the third step follows from Fact A.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$. ∎
@@
Lemma G.9 (Adaptive Softmax ).
@@
  • Let the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y)\}_{j=1}^{l})$.
@@
  • Let $B=O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
@@
$|f(y)-Z(q)|\leq|\epsilon_{s}Z(q)|+B+O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}).$
@@
The proof follows from the same idea as the proof of Lemma F.8, except that we use Theorem G.7 and the Lipschitz in Lemma G.8. ∎
Algorithm 7 Adaptive query data structure
2:members
3:   ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$ $\triangleright$ Algorithm 3
4:end members
[…]
6:  $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
7:  for $i=1\to l$ do
[…]
9:  end for
10:end procedure
11:procedure DistanceQuery($y\in[0,R]^{d}$)
12:  $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
13:  $r\leftarrow 0^{l}$
14:  for $i=1\to l$ do
15:   $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{DistanceQuery}(y)$
16:  end for
17:  return Median of $r$
18:end procedure
19:end datastructure
Theorem G.10 (Adaptive query Softmax data structure, formal version of Theorem 5.3).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 7) that uses $O(lnr)$ spaces to solve weighted Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
[…]
@@
    • $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma G.9. We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$. It can be shown that:
@@
G.4 Proof of Main Result